OpenStudy (anonymous):

what is the slope between the points (1,5) (1,-3)

OpenStudy (anonymous):

\(\dfrac{rise}{run} = \dfrac{\text{change in y}}{\text{change in x}} = \dfrac{y_2-y_1}{x_2-x_1}\)